# Quantum Computation by Cooling

### v2

@@ Abstract
Adiabatic quantum computation is a paradigmatic model aiming to solve a computational problem by finding the many-body ground state encapsulating the solution. However, its use of an adiabatic evolution depending on the spectral gap of an intricate many-body Hamiltonian makes its analysis daunting. While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, an analysis of such a scheme is surprisingly missing. Here, we propose a specific Hamiltonian model for this purpose. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system toward its ground state. At the same time, the measurement of the discarded qubits hints at the energy level structure of the system as a return. We show that quantum computation based on this cooling procedure is equivalent in its computational power to the one based on quantum circuits. We then exemplify the scheme with a few illustrative use cases for combinatorial optimization problems. In the first example, the cooling is free from any local energy minima, reducing the scheme to Grover’s search algorithm with a few improvements. In the second example, the cooling suffers from abundant local energy minima. To circumvent this, we implant a mechanism in the Hamiltonian so that the population trapped in the local minima can tunnel out by high-order transitions. We support this idea with a numerical simulation for a particular combinatorial optimization problem. We also discuss its application to preparing quantum many-body ground states, arguing that the spectral gap is a crucial factor in determining the time scale of the cooling.
@@ I Introduction
In this work, much of our focus will be placed on CQC for combinatorial optimization problems, which have favorable features for CQC and allow for the analytical treatment. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some additional advantages that CQC brings. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begin to pose challenges. In this case, high-order transitions, which are required to get out of local energy minima, become the bottleneck. These transitions require a time exponentially large in the order, which can increase with the system size in general. We perform the numerical simulation for a particular combinatorial optimization problem to support this idea. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and more complicated to analyze. We argue that the spectral gap above the ground state is a crucial factor in determining the time scale of the cooling. This is a natural consequence of the energy-time uncertainty principle.
@@ II Notation and working principles
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle$ to $|z^{\prime}\rangle$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|\langle z^{\prime}|\lambda H_{T}|z\rangle|$, where the influence of the degeneracy is ignored. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|\langle z^{\prime}|\lambda H_{T}|z\rangle|$ and couple it to the transition. Here, the subscript $m$ is the index for different cavities. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to prevent heating.
@@
$\begin{split}H&=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}\\ &\quad+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].\end{split}$ (2)
@@
Initially, the system is prepared in state $2^{-N/2}\sum_{z=0}^{2^{N}-1}|z\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. A single cycle of cooling is performed by unitarily evolving the system according to Hamiltonian (2) for an appropriate duration of time, measuring the cavity states, and resetting every cavity to the vacuum state. Here, the measurement of the cavity photon is not mandatory, but it would reveal some information on the current state and the internal transition structure of the problem Hamiltonian.
The computation is done by repeating this cycle. The required number of cycles depends on the problem and strategy. One possible strategy is to repeat until the cavity photon is not detected for an enough number of cycles. This signals with high probability that the system has reached either the ground state or local minima. For NP problems that are guaranteed to have at least one solution, deciding which is the case is straightforward. Otherwise, one may repeat the same computation to increase the confidence. Given that the unknown ground state is not verifiable in general, the decision is probabilistic after all.
@@
III Equivalence to circuit models
@@
A similar equivalence relation can be shown for CQC. Again, showing that CQC can be simulated efficiently in a quantum circuit is straightforward. We thus focus on the other direction. Suppose a quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state right after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
@@
[…]
where the parameter $\Delta>0$ is an arbitrary energy scale, and the other parmeters are chosen as $M=1$, $\omega_{1}=\Delta$, and $\alpha_{0}=0$. If the initial state is chosen to be $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$, the entire transition dynamics is restricted to a subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C},0\leq t\leq T\}$, as depicted in Fig. 1. It can be seen that the average time needed to reach the ground state, which is the final state of the quantum circuit, is polynomial in $T$. As the transition rate $\lambda$ is restricted by the energy gap $\Delta$, the computation time is $\mathcal{O}(T/\Delta)$.
@@ IV.2 Reproduction of Grover’s search algorithm
$\begin{split}H&=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}\\ &\quad+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),\end{split}$ (5)
@@
In this basis, the off-diagonal element becomes ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. The computation time is thus $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. One advantage of CQC implementation is that the end of computation is signaled by a detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution states and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to a decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.
@@ IV.3 Alternative approach
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\rightarrow|z_{1}\rangle_{P}\rightarrow\cdots|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. The perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ are different, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions come from the cases wherein the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (simply put, the transition diagram has a mirror symmetry). This rule is easily verifiable with simple numerical simulations, which we do not present here.
@@
Now taking the cavity modes into account, the condition can also be met when an intermediate transition $|z_{i}\rangle_{P}\rightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\rightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnelling-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$ are expected to play an important role in overcoming local minima.
@@
To illustrate the above idea, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [38, 30]). Specifically, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each calculation step into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has a value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has a value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{2}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.

### v4

# Markovian Quantum Computation

@@ Abstract
Adiabatic quantum computation is a paradigmatic model aiming to solve a computational problem by finding the many-body ground state encapsulating the solution. However, its use of an adiabatic evolution depending on the spectral gap of an intricate many-body Hamiltonian makes its analysis daunting. While it is plausible to directly cool the final gapped system of the adiabatic evolution instead, the analysis of such a scheme on a general ground is missing. Here, we propose a specific Hamiltonian model for this purpose. The scheme is inspired by cavity cooling, involving the emulation of a zero-temperature reservoir. Repeated discarding of ancilla reservoir qubits extracts the entropy of the system, driving the system toward its ground state. At the same time, the measurement of the discarded qubits hints at the energy level structure of the system as a return. We show that quantum computation based on this cooling procedure is equivalent in its computational power to the one based on quantum circuits. We then exemplify the scheme with a few illustrative use cases for combinatorial optimization problems. In the first example, the cooling is free from any local energy minima, reducing the scheme to Grover’s search algorithm with a few improvements. In the second example, the cooling suffers from abundant local energy minima. To circumvent this, we implant a mechanism in the Hamiltonian so that the population trapped in the local minima can tunnel out by high-order transitions. We support this idea with a numerical simulation for a particular combinatorial optimization problem. We also discuss its application to preparing quantum many-body ground states, arguing that the spectral gap is a crucial factor in determining the time scale of the cooling.
@@ I Introduction
In this work, much of our focus will be placed on CQC for combinatorial optimization problems, which have favorable features for CQC and allow for the analytical treatment. We consider two extreme cases. The first case is free from local energy minima, but the transition rate is superpolynomially small in the system size. This case turns out to be equivalent to Grover’s quantum search algorithm, aside from some advantages that CQC brings. The second case is an opposite limit, wherein the transition rate is maximized, but local energy minima begin to pose challenges. The local minima can be overcome by introducing high-order tunnel-out transitions. We perform the numerical simulation for a particular combinatorial optimization problem to support this idea. Finally, we briefly discuss the case of quantum Hamiltonians. This case is heavily system-dependent and more complicated to analyze. We argue that the spectral gap above the ground state is a crucial factor in determining the time scale of the cooling. This is a natural consequence of the energy-time uncertainty principle.
@@ II Notation and working principles
If we consider the transition problem described by Hamiltonian $H_{P}+\lambda H_{T}$, the transition from $|z\rangle_{P}$ to $|z^{\prime}\rangle_{P}$ is allowed only when $|E(z)-E(z^{\prime})|\lesssim|{}_{P}\langle z^{\prime}|\lambda H_{T}|z\rangle_{P}|$, where the influence of the degeneracy is ignored. In order to ensure $E(z^{\prime})<E(z)$, we introduce cavity field $c_{m}$ with resonant frequency $\omega_{m}\gg|{}_{P}\langle z^{\prime}|\lambda H_{T}|z\rangle_{P}|$ and couple it to the transition. Here, the subscript $m$ is the index for different cavities. Let us denote by $|\cdot\rangle_{m}$ the photon number state of the cavity. If the cavity is initially empty, the transition from $|z\rangle_{P}|0\rangle_{m}$ to $|z^{\prime}\rangle_{P}|1\rangle_{m}$ is allowed only when $E(z^{\prime})\simeq E(z)-\omega_{m}$, providing a mechanism to prevent heating.
@@
$H=H_{P}+\sum_{m=1}^{M}\omega_{m}c_{m}^{\dagger}c_{m}+\lambda H_{T}\otimes\left[a_{0}I+\sum_{m=1}^{M}(c_{m}+c_{m}^{\dagger})\right].$ (2)
@@
Our cooling protocol is a stochastic process performed by repeating the following cycle. To simplify the description, suppose $a_{0}=0$ and the system is initially in state $|z_{0}\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$ with arbitrary $z_{0}$. The energy of this state is $E(z_{0})$. First, we let the system evolve unitarily under Hamiltonian (2). This causes the system to undergo coherent oscillations between the initial state and a superposition of all states $|z_{m}\rangle_{P}|1\rangle_{m}$ satisfying $|E(z_{0})-\{E(z_{m})+\omega_{m}\}|\lesssim|{}_{P}\langle z_{m}|\lambda H_{T}|z_{0}\rangle_{P}|$, where it is understood that $|z_{m}\rangle_{P}$ may represent a superposition of multiple energy levels of $H_{P}$ in cases of degeneracy. After a certain period (explained below), each cavity state is measured. If a cavity photon, say $|1\rangle_{m}$, is detected, the system collapses into state $|z_{m}\rangle_{P}|1\rangle_{m}$, where $|z_{m}\rangle_{P}$ has a lower energy than $|z_{0}\rangle_{P}$ by $\omega_{m}$. In this case, we empty the cavity by resetting the state to $|0\rangle_{m}$ so that the next cooling cycle starts from the lower-energy state $|z_{m}\rangle_{P}\bigotimes_{m=1}^{M}|0\rangle_{m}$. If no cavity photon is detected, the system simply collapses back into the initial state of the current cycle. Note that in any case, the cycle ends by emptying all cavities. This information-discarding process simulates a zero-temperature reservoir. During the cycle, heating is avoided as long as $\omega_{m}\gg|{}_{P}\langle z_{m}|\lambda H_{T}|z_{0}\rangle_{P}|$ for all $m$. Precise control of the operation time is not essential as the cooling cycle is repeated without the risk of heating; given that the probability of detecting a cavity photon in a single cycle is finite (depending on the duration), the expected number of cycles to lower the energy is also finite. Having said that, the optimal duration of the unitary evolution is given by the inverse of the oscillation frequency multiplied by $\pi/2$. However, as this optimal time varies for each transition path and multiple oscillations take place simultaneously, finding a single optimal duration is generally challenging. The case for $a_{0}=1$ is elaborated in Sec. IV.
The measurement of the cavity state allows us to monitor the progress of cooling. It is a reasonable strategy to repeat the cooling cycle until cavity photons are not detected for a sufficient number of cycles. This signals with high probability that the system has reached either the ground state or a local minimum. For NP problems that are guaranteed to have at least one solution, distinguishing between the two cases is straightforward. Otherwise, to increase confidence, one may repeat the computation, possibly varying the transition term $H_{T}$. Given that an unknown ground state is not verifiable in general, the decision is probabilistic after all.
@@
III Equivalence to the quantum circuit model
@@
A similar equivalence relation can be shown for CQC. Again, showing that CQC can be simulated efficiently in a quantum circuit is straightforward. We thus focus on the other direction: a quantum-circuit computation can be efficiently simulated in CQC. Suppose a quantum algorithm is run on a quantum circuit in $T$ time steps, where $U_{t}^{P}$ is applied to the qubits at the $t$-th time step. Let $|\phi_{t}\rangle_{P}\equiv U_{t}^{P}U_{t-1}^{P}\cdots U_{1}^{P}|\phi_{0}\rangle_{P}$ be the state right after applying $U_{t}^{P}$, and tag this state with the clock state as $|\phi_{t}\rangle_{P}|t\rangle_{C}$. The idea is to get $|\phi_{t}\rangle_{P}|t\rangle_{C}$ transformed into $|\phi_{t+1}\rangle_{P}|t+1\rangle_{C}$ in the course of an energy-lowering transition. This is enabled by choosing the terms in Hamiltonian (2) as
@@
$H_{T}$ $=$ $\sum_{t=0}^{T-1}(U_{t+1}^{P}\otimes|t+1\rangle_{C}\langle t|+\mathrm{H.c.}),$ (4)
where the parameter $\Delta>0$ is an arbitrary energy scale, and the other parmeters are chosen as $M=1$, $\omega_{1}=\Delta$, and $\alpha_{0}=0$. If the initial state is chosen to be $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$, the entire transition dynamics is restricted to a subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C},0\leq t\leq T\}$, as depicted in Fig. 1. Note that the off-diagonal element ${}_{P,C}\langle\phi_{t+1},t+1|\lambda H_{T}|\phi_{t},t\rangle_{P,C}=\lambda$ is constant for every $t$.
The cooling occurs in the same manner as the sideband cooling [34]. Initially, the system undergoes the Rabi oscillation between $|\phi_{0}\rangle_{P}|0\rangle_{C}|0\rangle_{1}$ and $|\phi_{1}\rangle_{P}|1\rangle_{C}|1\rangle_{1}$ with Rabi frequency $\lambda$. After a certain duration, the cavity state is measured. In case of measuring $|1\rangle_{1}$, we reset the cavity state to $|0\rangle_{1}$. The subsequent evolution then becomes a Rabi oscillation between $|\phi_{1}\rangle_{P}|1\rangle_{C}|0\rangle_{1}$ and $|\phi_{2}\rangle_{P}|2\rangle_{C}|1\rangle_{1}$. If $|0\rangle_{1}$ is measured, the same Rabi oscillation simply restarts. Repeating this cycle thus makes the state cascade down the energy levels. The optimal duration of the Rabi oscillation is $\pi/2\lambda$, for which the cavity photon is detected with unit probability. However, precise timing is not essential, as mentioned in the previous section. If the cavity photon is detected $n$ times, the current state of the system is $|\phi_{n}\rangle_{P}$. Consequently, the average time required to reach the final state is polynomial in $T$. As the Rabi frequency of the oscillation $\lambda\ll\Delta$ is constant at every cycle, the computation time is $\mathcal{O}(T/\Delta)$.
The above Rabi oscillation relies on the conventional rotating wave approximation, which neglects terms that do not preserve the energy of the unperturbed Hamiltonian $H_{P}$. The omitted term in this approximation induces heating albeit with a small probability. However, this heating is essentially harmless because it does not take the state out of the subspace spanned by $\{|\phi_{t}\rangle_{P}|t\rangle_{C}\}$. Even if the heating occurs (with a small probability), the mere effect is to regress the cooling by a single step, causing the transition $|\phi_{t}\rangle_{P}|t\rangle_{C}\rightarrow|\phi_{t-1}\rangle_{P}|t-1\rangle_{C}$. Consequently, the overall computation time of $\mathcal{O}(T/\Delta)$ remains rigorously correct.
As mentioned in the introduction, the equivalence of CQC to the quantum circuit model should be interpreted carefully. What we have shown above is that computations in one model can be converted to those in the other model up to polynomial overhead, indicating that they are computationally equivalent. However, it is important to note that the two models are motivated differently. CQC is basically a scheme to prepare a many-body ground state. While CQC is advantageous in cases where finding a ground state naturally accomplishes a computation, such as in combinatorial optimization, or finding a ground state itself is the objective, there is no inherent reason to prefer CQC over the quantum circuit model for ordinary quantum algorithms. The same argument applies to AQC as it shared the same motivation with CQC [19].
@@ IV.2 Reproduction of Grover’s search algorithm
$H=\sum_{z=0}^{2^{N}-1}\Delta f(z)|z\rangle_{P}\langle z|+\Delta c_{1}^{\dagger}c_{1}+\lambda\bigotimes_{i=1}^{N}\frac{I_{i}+X_{i}}{2}\otimes(c_{1}+c_{1}^{\dagger}),$ (5)
@@
The cooling process is thus very simple. The unitary evolution of the cooling cycle is a Rabi oscillation between $|\phi_{0}\rangle_{P}|0\rangle_{1}$ and $|\phi_{1}\rangle_{P}|1\rangle_{1}$ occurring with the Rabi frequency ${}_{P}\langle\phi_{0}|\lambda\bigotimes_{i}\frac{I_{i}+X_{i}}{2}|\phi_{1}\rangle_{P}\simeq\lambda n_{0}^{1/2}2^{-N/2}$. By repeating the cooling cycle until the cavity photon is detected, one can obtain the solution state $|\phi_{0}\rangle_{P}$. As the optimal duration of the Rabi oscillation is proportional to the inverse of the Rabi frequency, the computation time is given by $\mathcal{O}(\sqrt{2^{N}/n_{0}})$, identical to that of Grover’s algorithm. In comparison to Grover’s algorithm, one advantage of the CQC implementation is that the end of computation is signaled by the detection of the cavity photon. This is especially beneficial when the number of solutions $n_{0}$ is unknown.
In this example, the problem Hamiltonian has only two energy levels corresponding to the solution state and the rest. A naturally following question is if the performance can be improved when the problem Hamiltonian has more than two energy levels. It can be seen that with the transition term chosen above, more energy levels just make the computation slower. To see this, note that the quadratic speedup comes from the collective transition, the rate of which increases with the degeneracy of the energy level. The increase in the number of energy levels leads to a decrease in the degeneracy of each level, which diminishes the collective effect. Moreover, it results in an increased number of transitions required to reach the ground state.
@@ IV.3 Alternative approach
To elucidate this mechanism, consider an $n$-th order transition through the sequence $|z_{0}\rangle_{P}\leftrightarrow|z_{1}\rangle_{P}\leftrightarrow\cdots\leftrightarrow|z_{n}\rangle_{P}$. For the moment, suppose that $\alpha_{0}=1$ in Hamiltonian (2) and there is no cavity mode involved. For such a transition to occur, the first requirement is $E(z_{0})=E(z_{n})$. In this case, the $n$-th order perturbation theory yields the effective transition rate $\mathcal{O}[\lambda(\lambda/\Delta)^{n-1}]$. However, if the Stark shifts of $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ differ, the effective detuning breaks the condition $E(z_{0})=E(z_{n})$, suppressing the transition. Consequently, the most prominent transitions result from cases where the energy differences satisfy the condition $E(z_{i+1})-E(z_{i})=E(z_{n-(i+1)})-E(z_{n-i})$ for all $i$ (see Fig. 2 for typical cases).
Figure 2: Demonstration of representative high-order transitions for $\Delta=10\lambda$, where $\Delta$ and $\lambda$ denote the energy gap and the Rabi frequency, respectively. The plots on the right show the population in $|z_{n}\rangle_{P}$ as a function of the rescaled time, starting from the initial state $|z_{0}\rangle_{P}$. The results for (a) $n=2,3,4,5,6$ and (b) $n=2,4,6,8,10$ are overlaid in the plots. The indistinguishability of the individual curves confirms the reliability of the perturbation treatment.
Among various high-order transition channels, we demonstrate two representative cases in Fig. 2. The Hamiltonian employed for the figure is
$H=\sum_{j=0}^{n}E(z_{j})|z_{j}\rangle_{P}\langle z_{j}|+\lambda\sum_{j=0}^{n-1}\left(|z_{j}\rangle_{P}\langle z_{j+1}|+|z_{j+1}\rangle_{P}\langle z_{j}|\right).$ (9)
In Fig. 2(a), we have chosen $E(z_{0})=E(z_{n})=0$, $E(z_{1})=E(z_{2})=\cdots=E(z_{n-1})=\Delta$, and $\lambda=\Delta/10$. In this case, the $n$-th-order perturbation theory states that the oscillation between $|z_{0}\rangle_{P}$ and $|z_{n}\rangle_{P}$ occurs with frequency $\Omega_{n}\simeq\Delta(\frac{\lambda}{\Delta})^{n}$. The plot on the right in Fig. 2(a) shows the population in $|z_{n}\rangle_{P}$ with respect to the rescaled time (in unit of $\pi/\Omega_{n}$) when the initial state is $|z_{0}\rangle_{P}$. The plot includes the curves for $n\in\{2,3,4,5,6\}$ and they are almost indistinguishable, confirming the reliability of the perturbation treatment. In Fig. 2(b), we have chosen $E(z_{j})=j\Delta$ for $j\leq n/2$ and $E(z_{j})=(n-j)\Delta$ for $j\geq n/2$. In this case, the oscillation frequency is given by $\Omega_{n}\simeq\Delta(\frac{\lambda}{\Delta}\frac{\lambda}{2\Delta}\cdots\frac{\lambda}{(n/2-1)\Delta})^{2}\frac{\lambda}{(n/2)\Delta}$ from the perturbation theory. The plot on the right shows the population in $|z_{n}\rangle_{P}$ for the initial state $|z_{0}\rangle_{P}$. The curves for $n\in\{2,4,6,8,10\}$ are plotted with respect to the rescaled time and they are again almost indistinguishable.
@@
Now, let us take the cavity modes into account. Note that the time evolution remains identical when any intermediate transition $|z_{i}\rangle_{P}\leftrightarrow|z_{i+1}\rangle_{P}$ is replaced by $|z_{i}\rangle_{P}|0\rangle_{m}\leftrightarrow|z_{i+1}^{\prime}\rangle_{P}|1\rangle_{m}$ with $E(z_{i+1})=E(z_{i+1}^{\prime})+\omega_{m}$. This is always possible provided $E(z_{i+1}^{\prime})<E(z_{i+1})$ and the appropriate cavity mode exists. By incorporating multiple cavity modes, various tunnel-out transition channels can appear. For instance, the lambda-type-like transitions characterized by $E(z_{1})-E(z_{0})=E(z_{n-1})-E(z_{n})+\omega_{m}$ with $n\geq 2$, equivalent to the case in Fig. 2(a), are expected to play an important role in overcoming local minima.
@@
To illustrate this approach to the cooling, we consider a particular integer factoring algorithm (not to be confused with Shor’s algorithm [39, 30]). Specifically, we follow the integer multiplication procedure as in the elementary arithmetic and turn it into a combinatorial optimization problem. The problem Hamiltonian is constructed as follows. Consider a multiplication of two 3-bit integers $\bar{x}\equiv(x_{2}x_{1}x_{0})_{2}$ and $\bar{y}\equiv(y_{2}y_{1}y_{0})_{2}$ resulting in a 6-bit integer $\bar{z}\equiv(z_{5}z_{4}\cdots z_{0})_{2}$, incorporating additional four carry bits $\bar{c}\equiv(c_{3}c_{2}c_{1}c_{0})_{2}$. For a given value of $\bar{z}$, we turn each calculation step into an energy term in the problem Hamiltonian. Here, 10 qubits are needed to encode $\bar{x}$, $\bar{y}$, and $\bar{c}$, while $\bar{z}$ is hard-coded in the problem Hamiltonian. For example, the first energy term has a value zero if $x_{0}y_{0}=z_{0}$ and $\Delta$ otherwise, the second has a value zero if $x_{1}y_{0}+x_{0}y_{1}=(c_{0}z_{1})_{2}$ and $\Delta$ otherwise, and so on. The ground state, representing the solution of the factoring, has energy zero as it satisfies all the conditions, and has a two-fold degeneracy as $\bar{x}$ and $\bar{y}$ are interchangeable.